# Large Language Model Evaluation via Matrix Nuclear-Norm

### v1

@@ A.3 Analysis of Algorithmic Complexity
A.3.1 Time Complexity Analysis
@@
A.3.2 Experimental Validation and Comparative Analysis
@@
A.3.3 Conclusion of the Complexity Analysis
@@
A.4 Proof of Theorem 1: Detailed Version
Here we present the proof of Theorem 1, demonstrating the strictly inverse monotonic relationship between $H(A)$ and $\|A\|_{F}$.
Given a matrix $A\in\mathbb{R}^{B\times C}$, where each element is non-negative and each row sums to 1, i.e.,
$\sum_{j=1}^{C}A_{i,j}=1\quad\text{for}\ i=1,2,\ldots,B.$
We define the Shannon entropy $H(A)$ as:
$H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log(A_{i,j}),$
and the Frobenius norm $\|A\|_{F}$ as:
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}|A_{i,j}|^{2}}.$
Step 1: Analysis for a Fixed Row Vector
For a fixed row $i$, let $a_{j}=A_{i,j}$ such that $a_{j}\geq 0$ and $\sum_{j=1}^{C}a_{j}=1$. The entropy for this single row is given by:
$H_{i}=-\sum_{j=1}^{C}a_{j}\log a_{j},$
and the Frobenius norm for this row vector $a$ is:
$\|a\|_{2}=\sqrt{\sum_{j=1}^{C}a_{j}^{2}}.$
To determine when $H_{i}$ is maximized or minimized, we apply the method of Lagrange multipliers. Consider the function:
$L(a_{1},a_{2},\ldots,a_{C},\lambda)=-\sum_{j=1}^{C}a_{j}\log a_{j}+\lambda\left(\sum_{j=1}^{C}a_{j}-1\right).$
Taking partial derivatives with respect to $a_{j}$ and setting them to zero:
@@ A.4 Proof of Theorem 1: Detailed Version
Since $\sum_{j=1}^{C}a_{j}=1$, we find $a_{j}=\frac{1}{C}$ for all $j$. Thus, $H_{i}$ is maximized when all $a_{j}$ are equal. In this case,
$\|a\|_{2}=\sqrt{\sum_{j=1}^{C}\left(\frac{1}{C}\right)^{2}}=\sqrt{\frac{1}{C}}.$
Conversely, $H_{i}$ reaches its minimum when one element $a_{k}=1$ and all other $a_{j}=0$. At this point,
$\|a\|_{2}=\sqrt{1^{2}+0}=1.$
This demonstrates that $H_{i}$ and $\|a\|_{2}$ are inversely related for each row.
Step 2: Inverse Monotonicity for the Entire Matrix
Since $H(A)$ is the average of all $H_{i}$ values:
$H(A)=\frac{1}{B}\sum_{i=1}^{B}H_{i},$
and $\|A\|_{F}$ is the square root of the sum of the squared norms $\|a\|_{2}^{2}$ for each row:
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\|a\|_{2}^{2}},$
the inverse relationship between $H_{i}$ and $\|a\|_{2}$ implies that $H(A)$ and $\|A\|_{F}$ are also inversely related across the entire matrix $A$.
Step 3: Determining the Bounds of $\|A\|_{F}$
To find the bounds, first note that when $a_{j}=\frac{1}{C}$ for all $j$,
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}\left(\frac{1}{C}\right)^{2}}=\sqrt{B\cdot\frac{1}{C}}=\sqrt{\frac{B}{C}},$
which corresponds to the minimum discriminability.
For the upper bound, when each row has only one element equal to 1,
$\|A\|_{F}=\sqrt{\sum_{i=1}^{B}1}=\sqrt{B},$
indicating maximum discriminability.
Step 4: Practical Implications
The inverse monotonic relationship between $H(A)$ and $\|A\|_{F}$ is significant in machine learning and classification tasks, where maximizing $\|A\|_{F}$ implies achieving greater certainty in predictions. This property can be used to evaluate model performance, where higher $\|A\|_{F}$ corresponds to a model that is more confident and discriminative in its predictions. Therefore, $H(A)$ and $\|A\|_{F}$ are strictly inversely monotonic, and the proof is complete.
A.5 Proof of Theorem 3
Given that $\|A\|_{F}$ is near $\sqrt{B}$, we want to approximate the $j$-th largest singular value $\sigma_{j}$ as $\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right)$. This approximation will be derived by examining the contributions of the columns of $A$.
1. Decomposition of $A$ and Eigenvalues: By singular value decomposition (SVD), $A=U\Sigma V^{T}$, where $\Sigma$ is a diagonal matrix containing singular values $\sigma_{1},\sigma_{2},\ldots,\sigma_{D}$.
When looking at the product $A^{T}A$, the element at position $(j,j)$ is:
$(A^{T}A)_{j,j}=\sum_{i=1}^{B}A_{i,j}^{2}$
This sum represents the total squared magnitude of the entries in the $j$-th column of $A$.
2. Connecting to Singular Values: Since singular values represent the contribution of orthogonal projections of $A$, if $\|A\|_{F}$ is near $\sqrt{B}$, the main contribution to the nuclear norm $\|A\|_{*}$ comes from columns with the largest sums of squared values $\sum_{i=1}^{B}A_{i,j}^{2}$.
3. Approximation: For matrices with well-distributed entries in each column (a common property when $\|A\|_{F}\approx\sqrt{B}$), the top singular values can be effectively estimated by sorting these column norms. Hence, the largest singular values $\sigma_{j}$ approximately correspond to the largest column norms $\sqrt{\sum_{i=1}^{B}A_{i,j}^{2}}$.
Therefore, for each $j\in\{1,\ldots,D\}$,
$\sigma_{j}\approx\text{top}\left(\sum_{i=1}^{B}A_{i,j}^{2},j\right).$
This concludes the proof that the $j$-th largest singular value $\sigma_{j}$ can be approximated by the top $j$ values of column-wise norms of $A$, as stated in Theorem 3.

### v3

@@ A.4 Analysis of Algorithmic Complexity
A.4.1 Time Complexity Analysis
@@
A.4.2 Experimental Validation and Comparative Analysis
@@
A.4.3 Conclusion of the Complexity Analysis
@@
A.5 Proof of Theorem 1
We prove the strictly inverse monotonic relationship between the entropy $H(A)$ and the Frobenius norm $\|A\|_{F}$ for a non-negative matrix $A\in\mathbb{R}^{B\times C}$ where each row represents a probability distribution:
$\sum_{j=1}^{C}A_{i,j}=1,\quad A_{i,j}\geq 0,\quad\forall i=1,\ldots,B.$
Definitions:
  • Entropy:
    $H(A)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}\log(A_{i,j})$
  • Frobenius norm:
    $\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\sum_{j=1}^{C}A_{i,j}^{2}}$
Step 1: Single-Row Analysis
For a row $\mathbf{a}=[a_{1},\ldots,a_{C}]$ with $\sum_{j}a_{j}=1$:
  • Row entropy: $H_{i}=-\sum_{j=1}^{C}a_{j}\log a_{j}$
  • Row norm: $\|\mathbf{a}\|_{2}=\sqrt{\sum_{j=1}^{C}a_{j}^{2}}$
Extrema via Lagrange Multipliers:
The Lagrangian $L=-\sum_{j}a_{j}\log a_{j}+\lambda(\sum_{j}a_{j}-1)$ yields:
@@ A.5 Proof of Theorem 1
Normalization gives $a_{j}=\frac{1}{C}$, achieving:
  • Maximum entropy: $H_{i}=\log C$
  • Minimum norm: $\|\mathbf{a}\|_{2}=\sqrt{\frac{1}{C}}$
Minimum entropy occurs when $a_{k}=1$ (one-hot vector):
  • Minimum entropy: $H_{i}=0$
  • Maximum norm: $\|\mathbf{a}\|_{2}=1$
Monotonicity: For fixed $C$, $H_{i}$ and $\|\mathbf{a}\|_{2}$ are strictly inversely monotonic (shown via derivative analysis or majorization theory).
Step 2: Matrix-Level Generalization
For the full matrix:
  • $H(A)=\frac{1}{B}\sum_{i=1}^{B}H_{i}$
  • $\|A\|_{F}=\sqrt{\sum_{i=1}^{B}\|\mathbf{a}_{i}\|_{2}^{2}}$
Key Observation: If each row’s entropy $H_{i}$ decreases (increases), its norm $\|\mathbf{a}_{i}\|_{2}$ increases (decreases). Thus: - $\|A\|_{F}^{2}=\sum_{i=1}^{B}\|\mathbf{a}_{i}\|_{2}^{2}$ decreases (increases) as $H(A)$ increases (decreases).
Step 3: Norm Bounds
Maximum $\|A\|_{F}$: When all rows are one-hot:
$\|A\|_{F}=\sqrt{B}$
Minimum $\|A\|_{F}$: When all rows are uniform:
$\|A\|_{F}=\sqrt{\frac{B}{C}}$
Step 4: Implications for LLMs
The inverse monotonicity implies:
  • High $\|A\|_{F}$: Concentrated predictions (low entropy, high confidence).
  • Low $\|A\|_{F}$: Dispersed predictions (high entropy, high diversity).
Thus, $\|A\|_{F}$ serves as a proxy for evaluating LLM confidence-diversity tradeoffs.
Conclusion
The strict inverse monotonicity between $H(A)$ and $\|A\|_{F}$ is rigorously established, justifying $\|A\|_{F}$ as a metric for LLM evaluation.
A.6 Proof of Theorem 3
Assuming $\|A\|_{F}\approx\sqrt{B}$ and the columns of $A$ are approximately orthogonal, we approximate the $j$-th largest singular value $\sigma_{j}$ as the $j$-th largest column norm of $A$. Formally,
$\sigma_{j}\approx\text{top}\left(\sqrt{\sum_{i=1}^{B}A_{i,j}^{2}},\ j\right),$
where $\text{top}(S,j)$ denotes the $j$-th largest element in set $S$. This approximation holds under the following analysis:
1. Decomposition and Gram Matrix: Let $A=U\Sigma V^{T}$ be the SVD of $A$, where $\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{D})$ with $D=\min(B,C)$. The diagonal entries of the Gram matrix $A^{T}A$ are:
$(A^{T}A)_{j,j}=\sum_{i=1}^{B}A_{i,j}^{2}=\|\mathbf{a}_{j}\|_{2}^{2},$
where $\mathbf{a}_{j}$ is the $j$-th column of $A$.
2. Relating Column Norms to Singular Values: When columns of $A$ are nearly orthogonal, $\sigma_{j}\approx\|\mathbf{a}_{j}\|_{2}$. Under $\|A\|_{F}\approx\sqrt{B}$, the nuclear norm $\|A\|_{\star}=\sum_{j=1}^{D}\sigma_{j}$ is dominated by the largest column norms.
3. Singular Value Approximation: For matrices with low column-wise correlations, the $j$-th singular value satisfies:
$\sigma_{j}\approx\text{top}\left(\{\|\mathbf{a}_{k}\|_{2}\mid 1\leq k\leq C\},\ j\right).$
4. Efficient Nuclear Norm Approximation: The batch nuclear norm is approximated as:
$\|\hat{A}\|_{\star}=\sum_{j=1}^{D}\text{top}\left(\{\|\mathbf{a}_{k}\|_{2}\},\ j\right).$
This approximation is valid when $A$ has approximately orthogonal columns, a condition implied by $\|A\|_{F}\approx\sqrt{B}$.